### v2

# On MCMC mixing for predictive inference under unidentified transformation models

Abstract
Reliable Bayesian predictive inference has long been an open problem under unidentified transformation models, since the Markov Chain Monte Carlo (MCMC) chains of posterior predictive distribution (PPD) values are generally poorly mixed. We address the poorly mixed PPD value chains under unidentified transformation models through an adaptive scheme for prior adjustment. Specifically, we originate a conception of sufficient informativeness, which explicitly quantifies the information level provided by nonparametric priors, and assesses MCMC mixing by comparison with the within-chain MCMC variance. We formulate the prior information level by a set of hyperparameters induced from the nonparametric prior elicitation with an analytic expression, which is guaranteed by asymptotic theory for the posterior variance under unidentified transformation models. The analytic prior information level consequently drives a hyperparameter tuning procedure to achieve MCMC mixing. The proposed method is general enough to cover various data domains through a multiplicative error working model. Comprehensive simulations and real-world data analysis demonstrate that our method successfully achieves MCMC mixing and outperforms state-of-the-art competitors in predictive capability.
Keywords: Bayesian nonparametrics; Identifiability; MCMC mixing; Predictive inference; Prior information level.
1 Introduction
We study the linear transformation model (Cuzick,, 1988),
$h(y)=\bm{\beta}^{\text{T}}\bm{z}+\epsilon,$ (1)
where $y\in\mathcal{Y}\subset\mathbb{R}$ is the response, $\bm{z}\in\mathcal{Z}\subset\mathbb{R}^{p}$ is the $p$-dimensional vector of covariates, $\bm{\beta}\in\mathbb{R}^{p}$ is the corresponding vector of regression coefficients, $h(\cdot)$ is a strictly increasing function, and $\epsilon$ is the continuous error term with cumulative distribution function (CDF) $F_{\epsilon}$. Over the past decades, numerous studies have contributed to statistical inference under the transformation model (1) (Horowitz,, 1996; Linton et al.,, 2008; Hothorn et al.,, 2018; Kowal and Wu,, 2024, among others), and may be categorized into two approaches.
i) A common strategy is the semiparametric regression approach that imposes a transformation on a specified reference distribution for the model error $\epsilon$ (Chen et al.,, 2002; Hothorn et al.,, 2014; Siegfried et al.,, 2023; Carlan et al.,, 2024; Kowal and Wu,, 2024; Brachem et al.,, 2024, among others). This strategy is straightforward and readily implementable, though it may encounter the risk of model misspecification. With this reference distribution strategy, most consistency results were established under log-concave-like assumptions on $F_{\epsilon}$ (Zeng and Lin,, 2006; Hothorn et al.,, 2018, among others). Such assumptions may fail in practice; in say, a normal mixture regression scenario (Soffritti and Galimberti,, 2011; Kasahara and Shimotsu,, 2015).
ii) A second approach is to allow $h$ and $F_{\epsilon}$ to both be unspecified in model (1). Identification conditions such as scale and location normalization constraints were imposed on either $h$ or $F_{\epsilon}$ (Horowitz,, 1996; Ye and Duan,, 1997; Chiappori et al.,, 2015). Despite their robustness and ideal theoretical properties, these methods are usually computationally intractable due to the employed kernel smoothing techniques. Although Chen, (2002) proposed a rank estimator of $h$ that does not require smoothing, he did not consider estimating $F_{\epsilon}$, and therefore cannot estimate the predictive distribution for future data. Mallick and Walker, (2003) imposed a constrained Polya tree prior for $F_{\epsilon}$ to identify model (1), but the posterior computation may not be stable since the posterior could suffer from slow mixing with an inappropriate center distribution (Müller et al.,, 2015).
In this article, we consider a third approach, where we allow the unspecified infinite-dimensional parameters $h$ and $F_{\epsilon}$ to be unidentified. That is, given the data, the likelihood is equal for a range of (infinite-dimensional) parameters $(h,\bm{\beta},F_{\epsilon})$; refer to Horowitz, (1996, pp. 105) for explicit description. We attempt to avoid complicated identification constraints for feasible computation. Specifically, we focus on Bayesian predictive inference (BPI), i.e. estimating the posterior predictive distribution (PPD) for future observations.
Though the BPI under unidentified models is conceptually doable, the key challenge that remains unresolved is the poor mixing of PPD value chains due to unidentifiability. Given $n$ observed data pairs $\mathcal{D}=\{y_{i},\bm{z}_{i}\}_{i=1}^{n}$ , suppose the future response $y^{*}$ is independent of $\mathcal{D}$ given the future covariates $\bm{z}^{*}$. The PPD value at a point $s\in\mathcal{Y}$ is $F_{y^{*}|\bm{z}^{*}}(s|\mathcal{D})=\int F_{y}(s|\bm{z}^{*},h,\bm{\beta},F_{\epsilon})d\pi(h,\bm{\beta},F_{\epsilon}|\mathcal{D}),$ where $F_{y}(s|\bm{z}^{*},h,\bm{\beta},F_{\epsilon})$ is the conditional CDF at $y=s$ under model (1) given parameters $(h,\bm{\beta},F_{\epsilon})$, and $\pi(h,\bm{\beta},F_{\epsilon}|\mathcal{D})$ is the joint posterior distribution. In practice, the PPD value is numerically approximated by Markov Chain Monte Carlo (MCMC) draws. Suppose that one draws $M$ parallel MCMC chains of the same length $N_{d}$, obtaining $MN_{d}$ draws of $(h^{(ml)},\bm{\beta}^{(ml)},F_{\epsilon}^{(ml)})$, for $m=1,\ldots,M$, $l=1,\ldots,N_{d}$. Then the PPD value at $s$ is approximated as the average of the PPD value chains:
$F_{y^{*}|\bm{z}^{*}}(s|\mathcal{D})\approx(MN_{d})^{-1}\sum_{m=1}^{M}\sum_{l=1}^{N_{d}}F_{y}(s|\bm{z}^{*},h^{(ml)},\bm{\beta}^{(ml)},F_{\epsilon}^{(ml)}).$
However, under the unidentified model (1), this approximation will NOT be reliable if the PPD value chains $F(s|\bm{z}^{*},h^{(ml)},\bm{\beta}^{(ml)},F_{\epsilon}^{(ml)})$ are poorly mixed in the sense that the M-chain PPD value samples do NOT converge to the stationary distribution. Poorly mixed PPD value chains usually incur poor BPI; see the lower expected log predictive densities (Yao et al.,, 2022, Corollary 5) for illustration.
Our solution is an adaptive scheme that leverages prior adjustment to achieve MCMC mixing. Our scheme operates like a bridge, on the one side is a new insight that under unidentified transformation models, the posterior variance is (asymptotically) dominated by the information level of the elicited Bayesian nonparametric priors (BNPs). This insight comes from a new asymptotic posterior variance decomposition, where the remainder term vanishes at a rate of $n^{-1}$, and the dominating term is fully determined by the hyperparameters in BNP elicitation; refer to Theorem 2. On the other side is the common principle that MCMC mixing occurs if the within-chain MCMC variance is sufficiently close to the posterior variance (Brooks et al.,, 2011, Section 6.1).
The insight and the principle motivate us to conceptualize a sufficient informativeness criterion: if the within-chain MCMC variance exceeds the dominating term (or its approximation) of the posterior variance, then the BNPs are sufficiently informative to reach MCMC mixing; accordingly, the prior information level is defined by the inverse of the dominating term. This criterion distinguishes the popular practice of computing the empirical between- and within-chain variances (Gelman and Rubin,, 1992; Brooks and Gelman,, 1998, among others) for discrimination of mixing only, since the analytic expression of the (approximated) prior information level (refer to Eq. (9)) can further activate expedient prior adjustment to achieve MCMC mixing; refer to Algorithm 1.
To derive the prior information level, we design an ideal BNP elicitation: a monotone spline model (Ramsay,, 1988) that possesses Lévy properties (Doksum,, 1974), and a Dirichlet process mixture model (Lo,, 1984) with a Weibull kernel. The hyperparameters induced from the hyperpriors for the BNP elicitation yield a neat analytic expression for an upper approximation to the prior information level. Consequently, prior adjustment is straightforwardly conducted by an adaptive hyperparameter tuning procedure without specific requirements on the initial values (refer to Section 5.1 for illustration).
The major contributions of this article are summarized as follows.
  • We contribute a robust and computationally feasible method for predictive inference under transformation models. Our methodology and theoretical results are general enough to cover the response types considered by conditional transformation models (Hothorn et al.,, 2014; Carlan et al.,, 2024). Specifically, we might be the first to establish the posterior inference theory under an unidentified nonparametric model, including the asymptotic posterior variance (Theorem 2) and the properness of the joint posterior (Theorem 3).
  • We contribute an easily implemented method to address the poor mixing of PPD value chains under unidentified transformation models. The hyperparameter tuning procedure is implemented under a general MCMC sampler Stan (Carpenter et al.,, 2017), releasing us from developing tricky samplers for multimodal target distributions (e.g. Pompe et al.,, 2020).
  • We contribute a quantile-knot I-spines BNP for nonnegative monotonic smooth functions. The proposed I-spline model enjoys lower model complexity (a few knots are enough) compared with other I-spline variants (e.g. Wang and Dunson,, 2011; Kim et al.,, 2017), while maintaining the root-$n$ posterior contraction rate that guarantees the asymptotic mixture of normals (Theorem 1).
  • We develop an R package BuLTM, for BPI under the transformation model (1). Comprehensive numerical studies demonstrate that BuLTM achieves the mixing of PPD value chains, and outperforms other state-of-art (SOTA) competitors in prediction tasks.
Organization. Section 2 presents an equivalent working model to model (1) and formulates the BNPs. Section 3 presents the adaptive scheme for prior adjustment to achieve MCMC mixing under the unidentified model. Section 4 establishes the properness of the joint posterior and introduces estimation of the parametric component. Simulations and applications to real-world data are presented in Sections 5 and 6 respectively. Section 7 contributes brief discussion. Technical proofs, additional simulation results, and other related details are collected in the Supplement. The companion R package BuLTM and the reproducible code for the numerical studies are available on GitHub https://github.com/LazyLaker/BuLTM.
2 Transformed modeling and nonparametirc priors
2.1 Multiplicative error working model
We first perform a transformation $\widetilde{h}$ on the response $y$ to transfer its support to (a subset of) $(0,\tau)$, for an arbitrary positive constant $\tau$. In this article, we consider the $\tau$-Sigmoid function $\widetilde{h}(y)=\tau/(1+e^{-y})$. Let $\widetilde{y}\in(0,\tau)$ be the transformed response and let $\circ$ be the composition of two functions operator. Based on model (1), we still have a transformation model $h^{*}(\widetilde{y})=\bm{\beta}^{T}\bm{z}+\epsilon$, where $h^{*}=h\circ\widetilde{h}^{-1}$ is n monotone increasing function, with $\widetilde{h}^{-1}$ being the (known) inverse function of $\widetilde{h}$.
Denote the transformation $\exp(h^{*})$ by $H$, and accordingly $\xi=\exp({\epsilon})$ with CDF $F_{\xi}$. We further have the following working model which is equivalent to model (1) in the sense of identical conditional distribution $F_{y|\bm{z}}$ under the two models
$H(\widetilde{y})=\xi\exp(\bm{\beta}^{\text{T}}\bm{z}).$ (2)
The equivalence is based on the fact that $Pr\{y\leq s|\bm{z}\}=Pr\{h(y)\leq h(s)|\bm{z}\}=Pr\{h^{*}(\widetilde{y})\leq h^{*}(s)|\bm{z}\}=Pr\{H(\widetilde{y})\leq H(s)|\bm{z}\}.$ Under working model (2), the following result holds naturally.
Proposition 1.
$H(0)=0$ if covariate $\bm{z}$ is independent of model error $\xi$.
The independence assumption between $\bm{z}$ and $\xi$ is general (Cuzick,, 1988; Horowitz,, 1996; Chen,, 2002). As a result, the space of $H$ is compressed to the space of nonnegative monotonic functions that passes through the origin.
Remark 1 (Nonlinearity).
The linear transformation model (1) and theworking model (2) are sufficiently general to incorporate nonlinear covariate effects. Let $\bm{z}=(z_{1},\ldots,z_{p})$. Let $\{\phi_{jk}\}_{k=1}^{K_{j}}$ be some basis functions (e.g. B-spline basis or Fourier basis) on $z_{j}$’s, for $j=1,\ldots,p$. Let $\widetilde{\bm{z}}_{j}=(\phi_{j1}(z_{j}),\ldots,\phi_{jK_{j}}(z_{j}))^{T}$ and $\otimes$ be the Kroneker product operator. Based on the tensor product basis (Pya and Wood,, 2015; Carlan et al.,, 2024), a smooth function $f:\mathbb{R}^{p}\to\mathbb{R}$ can be rewritten as $f(\bm{z})=\widetilde{\bm{\beta}}^{T}\widetilde{\bm{z}}$, where $\widetilde{\bm{z}}=\widetilde{\bm{z}_{1}}\otimes\ldots\otimes\widetilde{\bm{z}_{p}}$, and $\widetilde{\bm{\beta}}\in\mathbb{R}^{\prod_{j=1}^{p}K_{j}}$. To avoid the curse of dimensionality, one may consider an additive structure for $f$ (Linton et al.,, 2008; Chen et al.,, 2024) such that $f(\bm{z})=\sum_{j=1}^{p}f_{j}(z_{j})$, where $f_{j}(z_{j})=\sum_{k=1}^{K_{j}}\beta_{jk}\phi_{jk}(z_{j})\equiv\widetilde{\bm{\beta}}^{T}\widetilde{\bm{z}}$, where $\widetilde{\bm{z}}=(\widetilde{\bm{z}}_{1},\ldots,\widetilde{\bm{z}}_{p})^{T}$ and $\widetilde{\bm{\beta}}=(\beta_{11},\ldots,\beta_{1K_{1}},\ldots,\beta_{p1},\ldots,\beta_{pK_{p}})^{T}$.
2.2 Bayesian nonparametric priors
2.2.1 Quantile-knot I-splines prior
Given the data transformation, the observed data $\mathcal{D}$ become independent pairs of $\{\widetilde{y}_{i},\bm{z}_{i}\}_{i=1}^{n}$. For the transformed response $\widetilde{y}_{i}$ observed on the interval $D=(0,\tau)$, a natural method to model $H$ and its derivative $H^{\prime}$ is to use the monotone spline basis,
$H(s)=\sum_{j=1}^{K}\alpha_{j}B_{j}(s),\leavevmode\nobreak\ H^{\prime}(s)=\sum_{j=1}^{K}\alpha_{j}B_{j}^{\prime}(s),$ (3)
where $\{\alpha_{j}\}_{j=1}^{K}$ are positive coefficients to guarantee nondecreasing monotonicity, $\{B_{j}(s)\}_{j=1}^{K}$ are I-spline functions (Ramsay,, 1988) on $D$ and $\{B_{j}^{\prime}(s)\}_{j=1}^{K}$ are corresponding derivatives. Once $\{\alpha_{j}\}_{j=1}^{K}$ are specified, $H$ and $H^{\prime}$ are uniquely determined. By Proposition 1, we set $H(0)=0$ directly, unlike existing I-splines approaches that include an unknown intercept. A fundamental problem in spline modeling is how to specify the number of basis functions $K$, which is the sum of the number of interior knots and the order of smoothness $r$, defined by the existence of the (r1)(r-1)th order derivative. Empirically, the degree $r$ may take a value from $2$ to $4$ and we take the default value $r=4$ in R package splines2 (Wang and Yan,, 2021). The remaining task is to specify the number and locations of the interior knots.
We select interior knots from quantiles of the observed data, fitting a quantile-knots I-splines model, rather than using equally spaced knots. Let $\hat{F}_{n}(s)=n^{-1}\sum_{i=1}^{n}I(\widetilde{y}_{i}\leq s)$ be the empirical CDF of $\widetilde{y}$ and $\hat{Q}_{\widetilde{y}}(q)=\hat{F}_{n}^{-1}(q)=\inf\{s:q\leq\hat{F}_{n}(s)\}$ be the corresponding empirical quantile function, for $s\in(0,\tau)$ and $q\in(0,1)$. We first specify $N_{I}$, the number of interior knots (we set $N_{I}=4$ in this article as the default choice). Then the interior knots are set as $s_{j}=\hat{Q}_{\widetilde{y}}(j/N_{I})$, for $j=0,\ldots,N_{I}-1$. Such a quantile-knot configuration guarantees that the observed data lie uniformly between the knots.
Our quantile-knot I-spline BNP is appealing since one only needs a few knots rather than an increasing number of interior equally spaced knots (Wang and Dunson,, 2011; Kim et al.,, 2017), and hence has lower computational complexity. This BNP is not sensitive to the choice of the number of initial knots; refer to Supplement D.3. By assigning independent and identically distributed hyperpriors for the coefficients $\alpha_{j}$, the proposed quantile-knot I-spline BNP is closely related to the Lévy process (Doksum,, 1974); refer to Proposition A.1 in Supplement. This proposition guarantees the local asymptotics in Theorem 1 below.
2.2.2 Dirichlet process mixture model
For the prior for $F_{\xi}$ we consider the common Dirichlet process mixture (DPM) model (Lo,, 1984). Here we employ a truncated stick-breaking construction of the DPM, denoted as
$F_{\xi}(\cdot)=\int F_{0}(\cdot|\bm{u})dG(\bm{u}),\leavevmode\nobreak\ f_{\xi}(\cdot)=\int f_{0}(\cdot|\bm{u})dG(\bm{u}),\leavevmode\nobreak\ G=\sum_{l=1}^{L}p_{l}\delta_{\bm{u}_{l}},\leavevmode\nobreak\ \bm{u}_{l}\sim G_{0},$
where $F_{0}$ and $f_{0}$ are called kernels from a distribution family parameterized by $\bm{u}$, $L$ is a truncation number of the Dirichlet process, $p_{l}$ are corresponding sticking-breaking weights, and $\bm{u}_{l}$ are i.i.d. atoms from the base measure $G_{0}$. More justifications for the truncation level $L$ are deferred to Supplement B.
Note that $\xi$ is an arbitrary continuous positive random variable. In this article, we select the Weibull kernel for the DPM model,
$F_{\xi}(\cdot)=\sum_{l=1}^{L}p_{l}F_{w}(\cdot|\psi_{l},\nu_{l}),\leavevmode\nobreak\ f_{\xi}(\cdot)=(F_{\xi})^{\prime}=\sum_{l=1}^{L}p_{l}f_{w}(\cdot|\psi_{l},\nu_{l}),$ (4)
where $F_{w}(x|\psi_{l},\nu_{l})=1-\exp\{-(x/\psi_{l})^{\nu_{l}}\}$ and $f_{w}(x|\psi_{l},\nu_{l})=\nu_{l}\psi_{l}^{-\nu_{l}}x^{\nu_{l}-1}\exp\{-(x/\psi_{l})^{\nu_{l}}\}$ are the CDF and the pdf of the Weibull distribution with parameters $\{(\psi_{l},\nu_{l})\}_{l=1}^{L}$. Expression (4) yields a Weibull mixture model that has $L$ allocations of mixture components $(\psi_{l},\nu_{l})$, each with DP weights $p_{l}$.
The above Weibull kernel has at least two advantages: i) it can capture the shape of both monotone and nonmonotone hazards (Kottas,, 2006), and ii) it guarantees that the joint posterior under the unidentified working model (2) is proper; refer to Theorem 3.
2.2.3 Exponential hyperpriors and hyperparameters
Our nonparametric prior elicitation is completed by assigning hyperpriors to the parameters in the quantile-knot I-spline prior (3) and DPM model (4). Let $\bm{\alpha}=\{\alpha_{j}\}_{j=1}^{K}$, $\bm{p}=\{p_{l}\}_{l=1}^{L}$, $\bm{\psi}=\{\psi_{j}\}_{j=1}^{K}$, and $\bm{\nu}=\{\nu_{j}\}_{j=1}^{K}$. The hyperprior for $\bm{p}$ is naturally the stick-breaking prior (Sethuraman,, 1994). For $(\bm{\alpha},\bm{\psi},\bm{\nu})$, we assign exponential hyperpriors
$\pi(\bm{\alpha})=\prod_{j=1}^{K}\text{Exp}(\alpha_{j};\eta),\leavevmode\nobreak\ \pi(\bm{\psi})=\prod_{l=1}^{L}\text{Exp}(\psi_{l};\zeta),\leavevmode\nobreak\ \pi(\bm{\nu})=\prod_{l=1}^{L}\text{Exp}(\nu_{l};\rho).$ (5)
The rationale for employing the exponential hyperpriors is straightforward. For $\bm{\alpha}$, a Gamma hyperprior is preferable to link the I-splines model (3) with a Gamma process; refer to Proposition A.1 in the Supplement; for $(\bm{\psi},\bm{\nu})$ in the DPM with Weibull kernels, Gamma hyperpriors are becoming popular choices (Shi et al.,, 2019). We use exponential hyperpriors to avoid mathematically complicated formulations, though our theoretical results hold for arbitrary Gamma hyperpriors. With exponential hyperpriors, the nonparametric priors for $(H,F_{\xi})$ are parameterized by the hyperparameters $(\eta,\zeta,\rho)$.
3 Adaptive scheme for prior adjustment
The (infinite-dimensional) parameters in working model (2) are still unidentified. Suppose equation (2) holds for a special triplet solution $(H_{0},\bm{\beta}_{0},F_{\xi_{0}})$. Then equation (2) also holds on the set $\mathcal{C}\{(H,\bm{\beta},F_{\xi})\}=\{(c_{1}H_{0}^{c_{2}},c_{2}\bm{\beta}_{0},F_{c_{1}\xi_{0}^{c_{2}}})\}$ for any pair of positive constants $(c_{1},c_{2})\in\mathbb{R}_{+}^{2}$. In this section, we introduce an adaptive scheme to address the poorly mixing of PPD value chains under the unidentified working model (2). In Section 3.1, we focus on the asymptotic posterior variance first. If the posterior variance is divergent, no mixing results can be guaranteed; otherwise, it is possible for a general MCMC sampler to sufficiently explore the posterior uncertainty. In Section 3.2, we formulate the sufficient informativeness criterion based on the theoretical results in Section 3.1, and elucidate how to use the criterion to adaptively tune the hyperparameters to achieve MCMC mixing for trustworthy BPI. From now on, denote the expectation and variance operator with respect to a parameter $\theta$ under law $\pi(\theta)$ by $\mathbb{E}_{\theta}$ and $\mathbb{V}_{\theta}$ respectively, where $\pi(\theta)$ denotes the prior distribution of the parameter $\theta$.
3.1 Posterior variance under transformation models
Specifically, we focus on $\mathbb{V}\{H(s)|\mathcal{D}\}$, the posterior variance of $H(s)$ for some specific $s\in(0,\tau)$. Our motivation is the following conditional cumulative hazard function of the transformed response $\widetilde{y}$ given covariates $\bm{z}$. With the nonparametric prior elicitation (3) and (4), for $s\in(0,\tau)$, we have
[…]
In (6), the DPM components $(\bm{p},\bm{\psi},\bm{\nu})$ encounter the label-switching issue since the conditional cumulative hazard $\Lambda_{\widetilde{y}|\bm{z}}$ is invariant under any permutations of the indices of the allocation $(\psi_{l},\nu_{l})$; refer to (Mena and Walker,, 2015, pp. 1156) for general illustration. As a result, it is impossible to identify these parameters individually even if $F_{\xi}$ is specified. Fortunately, for a fixed $s\in(0,\tau)$, $H(s)$ does NOT encounter the label-switching issue as any permutations of DPM components has no impact on $\bm{\alpha}$ or $H(s)$, since $H$ is fully determined by $\bm{\alpha}$. This fact partially explains why we focus on the posterior variance of $H(s)$.
3.1.1 Preliminary: identified scenario
We start from a preliminary result in the case where $F_{\xi}$ is specified. With DPM model (4), specifying $F_{\xi}=F_{\xi_{0}}$ is equivalent to specifying $(\bm{p},\bm{\psi},\bm{\nu})$ at the ground truth $(\bm{p}_{0},\bm{\psi}_{0},\bm{\nu}_{0})$. The following conditions are further assumed.
(A1) All transformed response $\widetilde{y_{i}}$ are distinct.
(A2) There exists a constant $0<M_{\bm{z}}<\infty$ such that $\bm{z}||_{1}<M_{\bm{z}}$ with probability 1.
(A3) The prior $\pi(\bm{\beta})$ is continuous and $\pi(\bm{\beta})>0$ on $\mathbb{R}^{p}$.
(A4) The “true” $F_{\xi_{0}}$ can be expressed in the form of (4); in (4), $p_{l}>\delta$ for some positive constant $\delta$, $\sum_{l=1}^{L}\nu_{l}<\infty$ for $l=1,\ldots,L$.
Conditions (A1), (A2), and (A3) are general conditions in the literature for semiparametric Bernstein-von Mises (BvM) results (Kim,, 2006; Kim et al.,, 2017). Condition (A4) requires $F_{\xi 0}$ to be from the Weibull-kernel DPM family. In practice, (A4) can be relaxed so that the “true” $F_{\xi_{0}}$ falls into Weibull-kernel DPM’s Kullback-Leibler neighborhood, which is quite general (Wu and Ghosal,, 2008, Theorem 13).
The following theorem describes the asymptotic marginal posterior distribution of $H(s_{j})$, with “ground truth” $F_{\xi_{0}}$ given. The proof is deferred to Supplement A.
Theorem 1 (Asymptotic mixture of normals).
Suppose the “ground truth” $(\bm{p}_{0},\bm{\psi}_{0},\bm{\nu}_{0})$ is known. Let $H_{0}$ be the corresponding “true” transformation. Under conditions (A1) to (A4), with nonparametric priors (3) and (4), and hyperprior (5), for prespecified interior knots $s_{j}$ of the I-spline basis, for $j=1,\ldots,J$, as the data size $n\to\infty$, we have
$\pi[\sqrt{n}\{H(s_{j})-H_{0}(s_{j})\}|\mathcal{D},\bm{p}_{0},\bm{\psi}_{0},\bm{\nu}_{0}]\xrightarrow{\leavevmode\nobreak\ \leavevmode\nobreak\ \text{d}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ }\sum_{l=1}^{L}p_{l0}N\left\{0,\leavevmode\nobreak\ p_{l0}^{-1}\left(\frac{\psi_{l0}}{\nu_{l0}}\right)^{2}H_{0}(s_{j})^{\frac{2}{\nu_{l0}}-2}U_{l}(s_{j})\right\},$
where $U_{l}(s)=\int_{0}^{s}\{S_{l}^{0}(\mathcal{D},\bm{\beta}_{0})\}^{-1}d\Lambda_{l0}(s)$, with $\Lambda_{l0}(s)=\{H_{0}(s)/\psi_{l}\}^{\nu_{l}}$ and $S_{l}^{0}(\mathcal{D},\bm{\beta}_{0})$ is some positive constant depending on $\bm{\beta}_{0}$, $\nu_{l0}$ and data $\mathcal{D}$, for $l=1,\ldots,L$.
Theorem 1 relies on the fact that $H(s_{j})$ are sampled from a Lévy process $\mathcal{H}$ (refer to Proposition A.1 in the Supplement), which guarantees that $n^{1/2}(\mathcal{H}-H_{0})|\mathcal{D}$ weakly converges to a mixture of Gaussian processes. Consequently, the local posterior on a specific $s_{j}$ converges to a mixture of normals. The mixture of normals in Theorem 1 comes from the mixture structure of $F_{\xi 0}$. Theorem 1 also holds for censored data under the condition $\lim\limits_{n\to\infty}n_{1}/n>0$, where $n_{1}$ denotes the number of uncensored observations. In the special proportional hazard case where $L=1$ and $\psi_{10}=\nu_{10}=1$, Theorem 1 reduces to the BvM theorem (Kim,, 2006, Theorem 3.3).
Remark 2.
Theorem 1 can be extended to establish $\sqrt{n}$-consistency of $H(s)$ for all $s\in(0,\tau)$ by further assuming the following conditions: i) the number of knots $J\equiv J_{n}\to\infty$ as $n\to\infty$ such that $\max\limits_{j=1,\ldots,J}|s_{j}-s_{j+1}|\lesssim n^{-1/2}$; ii) $H_{0}$ is absolutely continuous on $0,\tau]$. Nevertheless, empirically a few knots are sufficient for estimation of PPDs. Meanwhile, to derive the sufficient informativeness criterion (refer to Criterion 1 in the next subsection), we only need the $\sqrt{n}$-consistency of $H$ with respect to each knot $s_{j}$.
3.1.2 Unidentified scenario
Under the unidentified model (2), where $F_{\xi}$ is drawn from the DPM model (4), the “ground truth” $F_{\xi_{0}}$ is no longer a fixed distribution, but, a sample of random functions. In this case, the posterior becomes multi-modal and the posterior variance will not vanish anymore. The following theorem formulates the asymptotic posterior variance of $H(s_{j})$ in the unidentified scenario.
Theorem 2 (Asymptotic posterior variance).
Assume conditions (A1) to (A4). Let $w_{j^{\prime}}=B_{j+j^{\prime}}(s_{j})-B_{j+j^{\prime}}(s_{j-1})$ for $j^{\prime}=1,\ldots,r$ in the I-splines model (3). As $n\to\infty$, under model (2), for parameter $\nu_{1}$ in DPM model (4), with hyperprior (5), there exist series of positive constants $\{c_{lj}\}_{l=1}^{L}$ and $\{r_{l}\}_{l=1}^{L}$ with $r_{1}=1$, such that for $j=1,\ldots,J$, $g_{s_{j}}(\nu_{1},\eta,\zeta)\equiv\zeta\sum_{l=1}^{L}c_{lj}^{\frac{1}{r_{l}\nu_{1}}}+\eta(j+\sum_{j^{\prime}=1}^{r}w_{j^{\prime}})^{-1},$ and
${\begin{aligned} \mathbb{V}\{H(s_{j})|\mathcal{D}\}&=\left[\mathbb{V}_{\nu_{1}}\left\{g_{s_{j}}^{-1}(\nu_{1},\eta,\zeta)\right\}+\mathbb{E}_{\nu_{1}}\left\{g_{s_{j}}^{-2}(\nu_{1},\eta,\zeta)\right\}\right]+O(n^{-1}),\\ &\equiv\mathcal{V}_{s_{j}}(\eta,\zeta,\rho)+O(n^{-1}).\end{aligned}}$ (7)
The first term on the RHS of (7) can be fully expressed in terms of the hyperparameters $(\eta,\zeta,\rho)$ since $\nu_{1}$ is integrated out. The second term is a remainder that vanishes at a rate of $n^{-1}$, which is a direct consequence of Theorem 1. Indeed, Theorem 2 is an explicit form of the following law of total variance under working model (2)
$\begin{aligned} \mathbb{V}\{H(s_{j})|\mathcal{D}\}=\underbrace{\mathbb{V}_{F_{\xi_{0}}}\{\mathbb{E}(H(s_{j})|\mathcal{D},F_{\xi_{0}})\}}_{\mathcal{V}_{s_{j}}(\eta,\zeta,\rho)}+\underbrace{\mathbb{E}_{F_{\xi_{0}}}\{\mathbb{V}(H(s_{j})|\mathcal{D},F_{\xi_{0}})\}}_{O(n^{-1})}.\end{aligned}$ (8)
In (8), we call $\mathbb{V}_{F_{\xi_{0}}}\{\mathbb{E}(H(s_{j})|\mathcal{D},F_{\xi_{0}})\}$ the mode variance since it is the variance of the posterior modes $\mathbb{E}(H(s_{j})|\mathcal{D},F_{\xi_{0}})$, and call $\mathbb{E}_{F_{\xi_{0}}}\{\mathbb{V}(H(s_{j})|\mathcal{D},F_{\xi_{0}})\}$ the local variance since it is the average of the variance around each local mode of the posterior. Obviously, if the model is identified, the mode variance disappears since the posterior mode is unique and fixed. In this unidentified scenario, the take-home messages of Theorem 2 are: i) under unidentified transformation models, asymptotically, the posterior variance will be dominated by the mode variance, which is expressed by the hyperparameters $(\eta,\zeta,\rho)$; ii) the large mode variance accounts for the poor mixing of MCMC chains, since the multiple chains should be sufficiently dispersed, if the single-chain variation is not enough to recover the mode variance.
To calculate the mode variance, we still needs to know the two positive constant series $\{c_{lj}\}_{l=1}^{L}$ and $\{r_{l}\}_{l=1}^{L}$, which are, however, unobservable. Fortunately, they have specific interpretations. Based on (6), given “true” $\bm{\nu}_{0}$, at each knot $s_{j}$, the ratio between “true” parameters $\bm{\psi}_{0}$ and $H_{0}$
$c_{lj}\equiv\left\{\frac{\psi_{l0}}{H_{0}(s_{j})}\right\}^{\nu_{l0}}=\left\{\frac{\psi_{l0}}{\sum_{j=1}^{K}\alpha_{0j}B_{j}(s_{j})}\right\}^{\nu_{l0}},\leavevmode\nobreak\ j=1,\ldots,K$
is uniquely determined, for $l=1,\ldots,L$. Furthermore, we can show that all “true” $\bm{\nu}_{0}$ fall in the space $\{(\nu_{10},\ldots,\nu_{L0}):\nu_{l0}/\nu_{10}=r_{l},l=2,\ldots,L\}$, where $r_{l}$ are some fixed positive constants; refer to Proposition A.4 in the Supplement. Based on these interpretations, we present an approximation to $\{c_{lj}\}_{l=1}^{L}$ in our adaptive scheme in the next subsection.
3.2 Sufficient informativeness for MCMC mixing
The above decomposition of the posterior variance motivates an adaptive scheme for prior adjustment to achieve the mixing of PPD value chains under unidentified transformation models. For a specific knot $s_{j}$, let $\mathbb{V}_{\text{WI}}(H(s_{j})|\mathcal{D})$ be the within-chain MCMC variance of $H(s_{j})$ of $M$ MCMC chains of length $N_{d}$:
$\mathbb{V}_{\text{WI}}(H(s_{j})|\mathcal{D})=M^{-1}\sum_{m=1}^{M}\sum_{l=1}^{N_{d}}\{H^{(ml)}(s_{j})-\bar{H}^{(m)}(s_{j})\}^{2},\leavevmode\nobreak\ \bar{H}^{(m)}(s_{j})=N_{d}^{-1}\sum_{l=1}^{N_{d}}H^{(ml)}(s_{j}).$
The following criterion assesses the MCMC mixing via the mode variance $\mathcal{V}_{s_{j}}$.
Criterion 1 (Sufficient informativeness criterion).
Under working model (2), the chains of PPD value of the transformed response $\widetilde{y}$ at the point $s_{j}$ are well mixed if $\mathbb{V}_{\text{WI}}\{H(s_{j})|\mathcal{D}\}\geq\mathcal{V}_{s_{j}}(\eta,\zeta,\rho),$ for $j=1,\ldots,J$. Then the BNPs for $H$ and $F_{\xi}$ are sufficiently informative.
Criterion 1 identifies the mixing of PPD value chains if the within-chain MCMC variance $\mathbb{V}_{\text{WI}}\{H(s_{j})|\mathcal{D}\}$ exceeds the mode variance. Criterion 1 stands on two facts: i) in well mixed MCMC chains, $\mathbb{V}_{\text{WI}}\{H(s_{j})|\mathcal{D}\}$ should approach (from below) the posterior variance $\mathbb{V}\{H(s_{j})|\mathcal{D}\}$ based on the ergodic theorem (Birkhoff,, 1942), and ii) the mode variance $\mathcal{V}_{s_{j}}(\eta,\zeta,\rho)$ is smaller than but dominates $\mathbb{V}\{H(s_{j})|\mathcal{D}\}$ based on Theorem 2. Consequently, we compare the within-chain variance with the mode variance to examine the convergence of MCMC chains to the target distribution.
Remark 3.
Note that the mode variance $\mathcal{V}_{s_{j}}$ is a function of the hyperparameters $(\eta,\zeta,\rho)$ for BNP elicitation (3) and (4). The hyperparameters $(\eta,\zeta,\rho)$ fully determine the uncertainties of the hyperpriors. Consequently, we call $\mathcal{V}_{s_{j}}^{-1}$ the “prior information level”: the smaller the mode variance, the more informative the BNPs are. Nevertheless, too informative BNPs can hinder the prior-to-posterior updating. Thus, our criterion uses the inverse of the within-chain MCMC variance $\mathbb{V}_{\text{WI}}^{-1}\{H(s_{j})|\mathcal{D}\}$ as a lower bound to determine the “sufficient” prior information level that achieves MCMC mixing and avoids slow posterior sampling.
Intuitively, by tuning the hyperparameters $(\eta,\zeta,\rho)$, we can increase the prior information level (or equivalently, decrease the mode variance $\mathcal{V}_{s_{j}}$) to satisfy Criterion 1. The remaining question is to approximate the unobservable constant series $\{c_{lj}\}_{l=1}^{L}$ and $\{r_{l}\}_{l=1}^{L}$.
By observing the form of $g_{s_{j}}(\nu_{1},\eta,\zeta)$ on the RHS of (7), if there exists a knot $s_{j}$ such that $c_{lj}\approx 1$, we can cancel $\nu_{1}$ and obtain a simple closed-form approximation for $\mathcal{V}_{s_{j}}(\eta,\zeta,\rho)$. Particularly, we only need a lower bound for $\mathcal{V}_{s_{j}}(\eta,\zeta,\rho)$ since Criterion 1 requires that the within-chain variance exceeds the information level. Therefore, to apply Criterion 1, it suffices to distinguish whether $c_{lj}<1$ or not. Based on (6), we have
$F_{\widetilde{y}|\bm{z}=\bm{0}_{p}}(s_{j})=1-\sum_{l=1}^{L}p_{l}\exp\left[-\left\{\frac{H(s_{j})}{\psi_{l}}\right\}^{\nu_{l}}\right]\equiv 1-\sum_{l=1}^{L}p_{l}\exp\left(-c_{lj}^{-\nu_{l}}\right).$
Suppose there exists $s_{j_{0}}$ such such that $F_{y^{*}|\bm{z}=\bm{0}_{p}}(s_{j})\geq 1-e^{-1}$. We have $\sum_{l=1}^{L}p_{l}c_{lj_{0}}^{-\nu_{l}}\geq 1$. That is, for $l=1,\ldots,L$, there exists at least one $c_{l_{j_{0}}}<1$. Then, we obtain the analytic expression of a lower approximation to $\mathcal{V}_{s_{j_{0}}}$ by replacing $c_{lj_{0}}$ to 1
${\mathcal{V}_{s_{j_{0}}}\geq\left(L\zeta+\frac{\eta}{j+\sum_{j^{\prime}=1}^{r}w_{j^{\prime}}}\right)^{-1}+\left(L\zeta+\frac{\eta}{j+\sum_{j^{\prime}=1}^{r}w_{j^{\prime}}}\right)^{-2}\equiv\widetilde{\mathcal{V}}_{s_{j_{0}}}(\eta,\zeta).}$ (9)
To use this approximation, we have to first specify the knot $s_{j}$. In practice, we consider the knot $s_{j_{0}}$ in the I-splines model (3) such that $s_{j_{0}}$ is the smallest among the knots that are greater than the $1-e^{-1}$ quantile of the transformed responses. We summarize this as the following criterion, an applicable version of Criterion 1.
Criterion 2 (Applicable sufficient informativeness criterion).
Under the working model (2), suppose we draw $M>1$ parallel MCMC chains. In the I-splines model (3), let $s_{j_{0}}=\hat{Q}_{\widetilde{y}}(q_{0}/N_{I})$ be the specific knot used for our criterion , where $q_{0}=\min\limits_{q=0,\ldots,N_{I}-1}\left\{1-q/N_{I}<e^{-1}\right\}$.
Then the BNPs are sufficiently informative if
$\mathbb{V}_{WI}\left\{H(s_{j_{0}})|\mathcal{D}\right\}\geq\widetilde{\mathcal{V}}_{s_{j_{0}}}(\eta,\zeta).$ (10)
Criterion 2 requires pre-configuration of the hyperparameter $\rho$ before MCMC sampling. We recommend specifying $\rho=1$ such that $E(\nu_{l})=1$, which is the same as the expectation of the LIO Weibull kernel hyperprior (Shi et al.,, 2019, pp. 690).
Algorithm 1 Adaptive tuning of hyperparameters $(\eta,\zeta)$ to reach MCMC mixing.
1:Specify $s_{j_{0}}$ in Criterion 2 based on data $\mathcal{D}$; set initial values for $(\eta,\zeta)\leftarrow(\eta_{0},\zeta_{0})$.
2:Draw $M>1$ MCMC chains with $N_{d}$ draws and examine the mixing by Criterion 2.
3:if inequality (10) does not hold then
4:  Select candidates $(\eta_{\text{new}},\zeta_{\text{new}})$ and set $\eta\leftarrow\eta_{\text{new}}$, $\zeta\leftarrow\zeta_{\text{new}}$; repeat 2 until Criterion 2 is met.
5:end if
The adaptive tuning procedure to select hyperparameters $(\eta,\zeta)$ is summarized in Algorithm 1. The choice of the initial values $(\eta_{0},\zeta_{0})$ is arbitrary. We recommend starting from very small values $(\eta_{0},\zeta_{0})$ to elicit “noninformative” BNPs. The selection of tuning candidates $\eta_{\text{new}}$ and $\zeta_{\text{new}}$ and the tuning procedure is illustrated and visualized in Section 5.1. The number of MCMC draws $N_{d}$ in each chain is related to the effective sample size and the MCMC sampler used. In Stan, we recommend using a chain length of $N_{d}=500$ (after a warm-up phase of the same length). Sensitivity analysis finds that longer MCMC chains will not change the tuning result. Detailed discussions on the chain length needed are deferred to Supplement D.1.
4 Joint posterior and parametric estimation
In this section, we attempt to answer the following two questions related to Bayesian inference under the working model (2): the first and most basic question is “is the joint posterior proper without identifiability?” and the next question is “can we estimate $\bm{\beta}$ with correct uncertainty quantification?”
4.1 The joint posterior is proper
Let $\bm{\theta}=(\bm{\alpha},\bm{\psi},\bm{\nu},\bm{p},\bm{\beta})$ be the collection of all parameters under working model (2). For the parametric component $\bm{\beta}$, we consider the objective improper uniform prior $\pi(\bm{\beta})\propto 1$. The following general conditions are assumed.
(B1) $\pi(\bm{p})$, $\pi(\bm{\psi})$, and $\pi(\bm{\nu})$ in model (4) and $\pi(\bm{\alpha})$ in model (3) are proper;
(B2) $0<K,L<\infty$ in models (4) and (3);
(B3) The kernel $f_{w}$ in model (4) satisfies $xf_{w}(x)<\infty$ for all $x>0$;
(B4) The $n\times p$ covariate matrix $\bm{Z}$ is of full rank $p$.
Conditions (B1) and (B2) are naturally satisfied, and (B3) is satisfied by the Weibull kernel. (B4) is similar to the condition $(ii)$ in de Castro et al., (2014), which is practical and easily validated. The following theorem tells us that, even with an improper prior for $\bm{\beta}$, the joint posterior of $\bm{\theta}$ is still proper. The proof is deferred to Supplement A.5.
Theorem 3.
Assume conditions (A1) to (A3) and (B1) to (B4). With the improper uniform prior for $\bm{\beta}$, under model (2), the posterior of $\bm{\theta}$ is proper.
Theorem 3 contradicts the results for unidentified parametric linear models, where proper priors lead to improper posteriors (Gelfand and Sahu,, 1999). This observation may imply that the infinite-dimensional parameters play a dominant role if a nonparametric model also has parametric components. Theorem 3 can be further extended to right-censored data by relaxing (B4) to $\bm{Z}^{*}$, the $n_{1}\times p$ covariate matrix of uncensored observations is of full rank $p$.
4.2 Parametric estimation with posterior projection
Under unidentified model (2), the marginal posterior intervals of $\bm{\beta}$ are generally too long to correctly quantify the uncertainty (Gelman et al.,, 2013). Therefore, we are driven to obtain the posterior of $\bm{\beta}^{*}$, the identified counterpart of $\bm{\beta}$ with certain normalization. Specifically, we consider unit-norm normalization such that $\bm{\beta}^{*}||_{2}=1$, where $||\cdot||_{2}$ denotes the Euclidian norm on $\mathbb{R}^{p}$. This differs from the element-one constraint, which needs extra effort to choose the covariate with coefficient fixed at 1 (Song et al., (2007); Lin et al., (2017); among others).
Rather than sampling $\bm{\beta}^{*}$ from the constrained space directly, we adopt the posterior projection (Sen et al.,, 2022) to project the marginal posterior of unconstrained $\bm{\beta}$ to the constrained parameter space of $\bm{\beta}^{*}$, the unit hyper-sphere $\text{St}(1,p)$ in $\mathbb{R}^{p}$. The metric projection operator $m_{\mathcal{A}}:\mathbb{R}^{p}\to\mathcal{A}$ of a set $\mathcal{A}$ is
[…]
5 Simulations
Extensive simulations are conducted to evaluate two aspects of the proposed method: i) how the proposed adaptive scheme guarantees the mixing of PPD value chains; and ii) how the proposed BuLTM package performs in predictive inference. In Section 5.1, we present examples that illustrate the hyperparameter tuning procedure for mixing PPD value chains; in Section 5.2, we compare BuLTM with other competitors.
Simulation setting. Our data setting covers two domains of response: (a) a real-valued response and (b) a positive response. In both settings, the simulated data are generated from model (1). In setting (a), the transformation $h$ is set to be the inverse (signed) Box-Cox function with $\lambda=0.5$, the same as the box-cox setting in Kowal and Wu, (2024). Two types of model error distributions are considered: (a.1) a standard normal distribution, a benchmark setting; (a.2) a normal mixture distribution, a model-misspecification setting for semiparametric methods. In setting (b), we allow the observations to be right-censored in a noninformative censoring scheme. Two types of model error distributions are considered: (b.1) an extreme-value distribution, the popular proportional hazard setting (Cox,, 1972); (b.2) a normal mixture distribution, a model-misspecification setting for semiparametric methods.
In each simulation, we generate $n=200$ samples as the training set and $n_{\text{test}}=20$ independent samples as the test set, and independently replicate the simulation runs $100$ times. An additional simulation setting (c) generated from nonlinear transformation models is deferred to Supplement C.
5.1 Visualization of the adaptive scheme for prior adjustment
We use examples from Setting (a.2) to illustrate the adaptive prior adjustment Algorithm 1 achieving well mixed PPD value chains. We examine two aspects of the mixing of PPD value chains: i) visualizing the trace plots of MCMC chains; ii) checking whether the rank normalized $\hat{R}$ statistic (Vehtari et al.,, 2021) exceeds $1.01$. We use the chains of the sum of the log posterior density of the observed data $\mathcal{D}$ given by MCMC samples, denoted by lp__ in Stan, as an alternative to the PPD value chains for simplicity. In each example, we use the $\tau$-Sigmoid transformation with $\tau=5$ as the data transformation mentioned in Section 2.1, and set the chain length for tuning as the default $N_{d}=500$. Examples in other settings and sensitivity analysis of the chain length are deferred to Supplement C and D respectively. In all examples, we set the hyperparameter $\rho=1$ as stated in Section 3.2.
Example 1.
Set initial values $(\eta_{0},\zeta_{0})=(0.01,0.01)$, yielding very vague priors for $H$ and $F_{\xi}$. After drawing MCMC samples, we compare the within-chain variance of $H(s_{j_{0}})$ with the inverse of the prior information level $\widetilde{\mathcal{V}}_{s_{j_{0}}}$. As shown in Figure 1, the within-chain MCMC variance is much less than $\widetilde{\mathcal{V}}_{s_{j_{0}}}(\eta_{0},\zeta_{0})$. Thus, we assert that the BNPs are not sufficiently informative to achieve mixing of PPD value chains. As evidence, Figure 1 shows that the MCMC traces of “lp__” are poorly mixed, with $\hat{R}=1.25$. Accordingly, the effective sample size (ESS) of lp__ is only 7, which is definitely insufficient.
Figure 1: (a) The curve of $\widetilde{\mathcal{V}}_{s_{j_{0}}}(\eta,\zeta)$ with $\eta=0.01$ fixed; horizontal line: the within-chain MCMC variance sampled with hyperparameters $(\eta,\zeta)=(0.01,0.01)$. (b) Trace plot of chains of lp__.
Figure 2: (a) The curve of $\widetilde{\mathcal{V}}_{s_{j_{0}}}(\eta,\zeta)$ with $\eta=0.01$ fixed; horizontal line: the within-chain MCMC variance sampled with hyperparameters $(\eta,\zeta)=(0.01,0.25)$. (b) Trace plot of chains of lp__.
Example 2.
Figure 1 illustrates hyperparameter tuning on $(\eta,\zeta)$. With $\eta=0.01$ fixed, candidates for updating $\zeta$ should enable the curve of $\widetilde{\mathcal{V}}_{s_{j_{0}}}$ against $\zeta$ to fall below the within-chain MCMC variance (the horizontal line). Meanwhile, the curve falls sharply on the interval $(0,.0.25]$, and decreases gently on the interval $[0,25,1]$. Consequently, we set $(\eta,\zeta)=(0,01,0.25)$ as the updated tuning hyperparameters. Figure 2 shows that the within-chain MCMC variance exceeds $\widetilde{\mathcal{V}}_{s_{j_{0}}}$, indicating that the BNPs are sufficiently informative. As a result, the MCMC chains of lp__ mix well as shown by Figure 2 with $\hat{R}=1.006$, demonstrating the efficacy of the tuning procedure. Meanwhile, the obtained ESS of 520 is sufficient to represent the log posterior densities.
According to Margossian and Gelman, (2023), reliability diagnostics ($\hat{R}$ and ESS) demonstrate that the estimated PPD in Example 2 is reliable. This example also illustrates that the hyperparameter configuration $(\eta,\zeta,\rho)=(0.01,0.25,1)$ achieves mixing of PPD value chains under setting (a.2). We further use this hyperparameter setting as the initial values throughout all numerical studies. Interestingly, this hyperparameter setting achieves mixing of PPD value chains in all our numerical studies.
We do NOT recommend increasing prior information by increasing $\eta$ only since in (9), $|\partial\widetilde{\mathcal{V}}_{s_{j_{0}}}/\partial\zeta|$ is much larger than $|\partial\widetilde{\mathcal{V}}_{s_{j_{0}}}/\partial\eta|$.
Figure 3: (a) The curve of $\widetilde{\mathcal{V}}_{s_{j_{0}}}(\eta,\zeta)$ with $\zeta=0.01$ fixed; horizontal line: the within-chain MCMC variance sampled with hyperparameters $(\eta,\zeta)=(5,0.01)$. (b) Trace plot of chains of lp__.
Example 3.
As a counter example, based on the initial values in Example 1, we fix $\zeta=0.01$ and set $\eta=5$, yielding a vague prior for $f_{\xi}$ and a highly informative prior for $H$. Unfortunately, this hyperparameter setting is insufficiently informative since the within-chain MCMC variance falls below $\widetilde{\mathcal{V}}_{s_{j_{0}}}$ as shown in Figure 3. Thus, the PPD value chains are poorly mixed as shown by Figure 3, with $\hat{R}=1.056$.
Note that it is meaningless to further increase $\eta$ since too informative a prior for $H$ leads to extremely slow sampling. This counter example illustrates another aspect of the utility of the proposed prior adjustment scheme: it strikes a balance between the noninformative priors that yield poor mixing and the too informative priors that hinder sampling.
5.2 Predictive capability evaluation
This subsection evaluates the predictive capability of the BuLTM package under transformation models. In BuLTM, we use the estimated PPD as the predictive distribution; for the predicted value, in Setting (a), we use the median of the estimated PPD; in Setting (B), we use the quantile of the estimated PPD that corresponds to the censoring rate.
Competitors. In Setting (a), competitors are the packages or open-source algorithms for fitting semiparametric transformation models. All competitors adopt the standard normal distribution as the reference distribution.
  • R package SeBR (Kowal and Wu,, 2024). We use the empirical CDF of predicted Monte Carlo samples as the predictive distribution, and use the default predicted value.
  • R code BCTM.lin (BCTM, Carlan et al.,, 2024). We use the default PPD as the predictive distribution and use predictive median as the predicted value.
  • Add-on R package tram (Siegfried et al.,, 2023). An add-on package in mlt (Hothorn,, 2020), the implementation of conditional transformation models (Hothorn et al.,, 2014, 2018). We use the default predictive distribution and the predicted meadian as the predicted value.
  • Python library liesel_ptm (PTM, Brachem et al.,, 2024). We use the default predictive distribution and predicted value.
In Setting (b), competitors are the packages for semiparametric survival models.
  • R package spBayesSurv (Zhou et al.,, 2020), a Bayesian package for semiparametric survival model fitting and model selection.
  • R package mlt (Hothorn,, 2020).
  • R package TransModel (Zhou et al.,, 2022), fitting a semiparametric transformation model based on Chen et al., (2002).
In Setting (b.1), all competitors use the correct reference distribution; in Setting (b.2), competitors use the reference distribution selected by spBayesSurv following the model selection procedure in Zhou and Hanson, (2018).
Assessments. We evaluate two capabilities: (i) the capability of recovering the predictive distribution; (ii) the performance of a single predicted value. Capability (i) is evaluated by the root integrated mean square error (RIMSE) between estimated predictive distribution $\hat{f}$ and the truth $f$: $\text{RIMSE}(\hat{f},f)=\sqrt{\int_{a}^{b}(\hat{f}(s)-f(s))^{2}ds}$ on an interval $(a,b)$. For capability (ii) assessment, we use the mean absolute error (MAE) in Setting (a), and the C-index (Harrell et al.,, 1982) in Setting (b).
Knot interpolation with censored data. We introduce a knot interpolation procedure to incorporate information from censored observations. Let $\widetilde{y}$ be the uncensored transformed observations, and $\widetilde{y}_{c}$ be the collection of both censored and uncensored observations. We begin with the $N_{I}$ interior knots specified by the quantiles of the uncensored observations, denoted by $s_{0},\ldots,s_{N_{i}-1}$. Then we interpolate the quantiles of $\widetilde{y}_{c}$ that are located far from the same quantiles of $\widetilde{y}$ and a complement of the knots. We summarize the two-step procedure in Algorithm 2. An example that illustrates and visualizes the procedure is deferred to Supplement E.
Algorithm 2 Knot interpolation with censored data
1:Configure initial knots. Let $N_{I}>1$ be the number of initial knots. For $j=0,\ldots,N_{I}-1$, let $s_{j}=\hat{Q}_{\widetilde{y}}(j/N_{I})$. Sort initial knots $0<s_{0}<\cdots<s_{N_{I}-1}<\tau$.
2:if  the transformed observations $\widetilde{y}$ are right-censored  then
3:  Let $\widetilde{y}_{c}$ be the collection of all observations, and $\widetilde{y}$ be the uncensored observations. For $s_{j}$ such that $|\hat{F}_{\widetilde{y}_{c}}(s_{j})-\hat{F}_{\widetilde{y}}(s_{j})|\geq 0.05$, interpolate a new knot $s_{j}^{*}=\hat{Q}_{\widetilde{y}_{c}}(j/N_{I})$.
4:  Output sorted series of $\{s_{1},\ldots,s_{j},s_{j}^{*},\ldots,s_{N_{I}-1}\}$ as final interior knots.
5:end if
Setting (a).
Figure 4: Box-plots of predictive assessments under Setting (a.1). (a), RIMSE; (b), MAE.
The box-plots of assessments among all replicative simulations in Settings (a.1) and (a.2) are presented in Figures 4 and 5 respectively. In setting (a.1), where all competitors correctly specify the model, BuLTM is competitive with SeBR in recovering predictive distributions (two-sided paired t-test $p$-value: 0.15 against SeBR), and significantly outperforms the remaining competitors except tram. However, BuLTM outperforms all competitors including tram in fitting the predicted values (one-sided paired t-test $p$-values : $1.147\times 10^{-5}$ against SeBR; $2.14\times 10^{-15}$ against BCTM). In setting (a.2), where all semiparametric methods encounter model misspecification, BuLTM significantly outperforms all competitors in both recovering predictive distributions and fitting the predicted values (one-sided paired t-test $p$-values : $0.0001$ against SeBR; $3.16\times 10^{-8}$ against BCTM).
Figure 5: Box-plots of predictive assessments under Setting (a.2). (a), RIMSE; (b), MAE.
Figure 6: Box-plots of predictive assessments under Setting (b.1). (a), RIMSE; (b), C-index.
Setting (b). The box-plots of assessments among all replicative simulations in Settings (b.1) and (b.2) are presented in Figures 6 and 7 respectively. In Setting (b.1), the commonly used proportional hazard setting, BuLTM significantly outperforms spBayesSurv (one-sided paired t-test $p$-values: $0.0002$) and TransModel ($2.03\times 10^{-5}$) in recovering the predictive distributions, and is comparable with spBayesSurv and TransModel in C-index, while mlt outperforms. In Setting (b.2), where the competitors encounter model misspecification, BuLTM significantly outperforms the competitors in recovering the predictive distributions, and slightly outperforms in C-index.
In summary, BuLTM is robust against model misspecification for both real-valued and positive responses, and is competitive (generally significantly outperforms) in both predictive distribution recovery and single-value predictions.
Figure 7: Box-plots of predictive assessments under Setting (b.2). (a), RIMSE; (b), C-index.
6 Applications
6.1 Auto MPG data
We first apply BuLTM to Auto MPG (Quinlan,, 1993), a benchmark machine learning dataset. The response is city-cycle fuel consumption in miles per gallon (MPG) and the predictors are 3 multivalued discrete and 5 continuous covariates. We preprocess the data by transforming all continuous predictors to $(0,1)$ and center the response to $\mathbb{R}$. We split the data into a 90% training set and a 10% test set and repeat the split for 10 runs to compare the out-of-sample predictive performance of BuLTM with other competitors. We allow for nonlinear covariate effects through covariate transformation with an additive structure such that $h(y)=\sum_{j=1}^{8}f_{j}(z_{j})+\epsilon$, where $f_{j}(z_{j})=\sum_{k=1}^{K}\beta_{jk}\phi_{j}(z_{j})$ and the $\phi_{j}$ are basis functions. On this dataset, we select the Fourier basis and set $K=8$. We apply the nonlinear model transformation to both BuLTM and SeBR, namely BuLTM.nonlin and SeBR.nonlin, respectively.
Two metrics on the test sets are used for assessment: i) the MAE between the predicted and the true values, and ii) the coverage probability (CP) of the 95% prediction intervals. We compare BuLTM with the competitors in Section 5 under simulation Setting (A). To reduce computational burden, we fit the linear model for BCTM without the smooth transformation. Thus, the results of BCTM can also be treated as a baseline model.
Figure 8: Box-plots of predictive assessments on the MPG dataset. (a), MAE; (b), CP, the horizontal line is the nominal level of coverage.
The box-plots of the assessment metrics are presented in Figure 8. We find that for both BuLTM and SeBR, the linear model enjoys lower MAE than the nonlinear model, indicating that a linear transformation model is adequate to fit the MPG data. On the MAE metric, BuLTM.lin is competitive with mlt (two-sided paired t-test $p$-value: 0.062) and significantly outperforms the other competitors. On the CP metric, both BuLTM and BuLTM.nonlin achieve the nominal coverage, while mlt fails to do so. This real-word example demonstrates the superiority of BuLTM in both fitting predicted values and recovering predictive distributions for real-valued data.
6.2 Heart failure clinical records data
The second real-world example is the heart failure clinical records data. The dataset records 299 heart failure patients collected at the Faisalabad Institute of Cardiology and at the Allied Hospital in Faisalabad, from April to December in 2015 (Ahmad et al.,, 2017). The dataset consists of 105 women and 194 men, with a range of ages between 40 and 95 years old. In the dataset, 96 subjects are recorded as dead and the remaining 203 are censored, leading to a censoring rate of 67.9%, which is relatively high. The dataset contains 11 covariates reflecting subject’s clinical, body, and lifestyle information. In this dataset, spBayesSurv selects the PH model and thus, TransModel specifies $r=0$, and mlt uses the reference distribution "MinExtrVal". We conduct 10 runs of 5-fold cross validation. The results of estimation of regression coefficients are deferred to Supplement F.1.
On the heart failure dataset, we use two metrics to assess the predictive capabilities: i) the C-index, where we use the $70\%$ quantile of the predictive distribution (close to the censoring rate) as the predicted survival time of a future observation; ii) the Integrated Brier Score (IBS Graf et al.,, 1999) on the follow-up time interval (the lower the IBS, the better the prediction).
Figure 9: Prediction comparison between BuLTM, spBayesSurv, and TransModel on the heart failure dataset; (a), C index; (b), Integrated Brier score; red dashed lines: the mean of the metrics.
Box-plots of the assessment metrics are presented in Figure 9. BuLTM significantly outperforms other competitors in both C-index and IBS: for C-index, the one-sided paired t-test $p$-values are $0.01$ against spBayesSurv, $0.0038$ against TransModel, and $0.0046$ against mlt; for IBS, the one-sided paired t-test $p$-values are $0.002$ against spBayesSurv, $1.47\times 10^{-5}$ against TransModel, and $2.8\times 10^{-4}$ against mlt. This example demonstrates the superiority of BuLTM in the prediction of censored data.
7 Discussion
Under unidentified transformation models, the proposed sufficient informativeness criterion extends the Gelman-Rubin (G-R) statistic (Gelman and Rubin,, 1992) from MCMC checking to covering prior adjustment, by taking another view of MCMC convergence. The G-R statistic diagnoses whether the MCMC transitions converge to the stationary distribution (Roy,, 2020). By contrast, we examine whether the within-chain MCMC variance is close to the true posterior variance (dominated by the inverse of prior information level; refer to Theorem 2) under unidentified transformation models.
As the AE and referees’ sharp insights have helped us clarify, the application scope of BuLTM covers general data domains through the multiplicative error working model. We want to emphasize that BuLTM offers a robust alternative toolbox to survival analysis in addition to spBayesSurv (Zhou and Hanson,, 2018) and TransModel (Zhou et al.,, 2022): it can estimate conditional survival functions and conditional hazard functions for future data, and provides a reliable Bayesian estimate of identified regression coefficient $\bm{\beta}$ with a tractable unit-norm constraint.
In our unidentified scenario caused by a flat likelihood, only a few discussions have mentioned that “weakly informative priors” may resolve the poor mixing phenomenon (Reich and Ghosh, (2019); McElreath, (2020)), but they did not quantify how “weak” the priors can be. In this sense, this article might be the first to quantitatively link prior informativeness with MCMC mixing under unidentified models: we rigorously demonstrate that the variance of multi-modal posteriors does not vanish with increasing data size, but rather is dominated by the prior information level; and we comprehensively illustrate how to achieve MCMC mixing by increasing the prior information through an analytic expression for prior information, an algorithm of hyperparameter tuning, and visualization of the whole procedure.
Our method addresses the poor mixing of PPD value chains under unidentified transformation models. Nonetheless, mixing of other parameters such as the Dirichlet process mixture components remains unsolved due to the label-switching issue. It is anticipated that an ordered Dirichlet process stick-breaking construction (Zarepour and Al Labadi,, 2012) can resolve the problem and speed up our MCMC sampling, but the implementation is so far unavailable in rstan since it needs the Boost C++ libraries. Our method may be further extended to other unidentified models such as latent Dirichlet allocation (LDA, Blei et al.,, 2003) and Bayesian additive regression trees (BART, Chipman et al.,, 2012), where new prior elicitation and new quantifications of prior information are needed.
SUPPLEMENTARY MATERIAL
Title:
Supplementary material for “On MCMC mixing for predictive inference under unidentified transformation models” (PDF file)